# How to Evaluate Monocular Depth Estimation?

### v1

@@ Abstract
Monocular depth estimation is an important task with rapid progress, but how to evaluate it remains an open question, as evidenced by a lack of standardization in existing literature and a large selection of evaluation metrics whose trade-offs and behaviors are not well understood. This paper contributes a novel, quantitative analysis of existing metrics in terms of their sensitivity to various types of perturbations of ground truth, emphasizing comparison to human judgment. Our analysis reveals that existing metrics are severely under-sensitive to curvature perturbation such as making flat surfaces wavy. To remedy this, we introduce a new metric based on relative surface normals, along with new depth visualization tools and a principled method to create composite metrics with better human alignment. Code and data are available at: https://github.com/princeton-vl/evalmde.
@@ 1 Introduction
However, how to evaluate monocular depth estimation remains an open problem. This is reflected by a lack of standardization in existing literature, and a bewilderingly large menu of evaluation metrics. It is in fact possible for two papers (e.g., (Bochkovskii et al., 2025; Wang et al., 2025a)) to report 5 to 7 different metrics each but with zero overlap [^1] Two metrics can share the same name but are different due to differences in depth alignment..
Even when models are compared under the same set of metrics, it is often not clear how to interpret the results. If a model performs well under one metric (say AbsRel after affine alignment on disparity) but poorly under another (say $\delta^{1}$ with no alignment), what does it mean? Often there lack clear answers, because the trade-offs between different metrics are not well understood.
The unusually large selection of metrics arises from a combination of decisions that need to be made when designing an evaluation metric. For example, the decisions can include how to compute error of individual depth values against ground truth (by difference in depth, log depth, or inverse depth), whether and how to discount error for faraway points, whether to binarize the error and if so with what threshold, and how to account for the unknown global scale through alignment (in depth or inverse depth), and whether to also allow an unknown global shift.
The complexity of these decisions is due to several issues inherent to monocular depth estimation that preclude straightforward comparisons of individual depth values. One is scale ambiguity. The global scale of a scene is fundamentally ambiguous: a scene can be a miniature replica of a larger identical scene. This scale ambiguity can also occur locally in a scene: it can be impossible to tell from a single image whether an airplane in the sky is a big one far away or small one close by. Even without any scale ambiguity, another issue is the unbounded range of possible depth values in the same scene—consider an ocean that extends to the horizon. In such cases, some form of normalization or weighting is necessary to prevent errors on faraway objects from dominating.
@@
In this paper we seek to improve our understanding of how to evaluate monocular depth estimation. Our goal is to study various evaluation metrics and shed light on their trade-offs and behaviors, and to develop a principled method to customize or combine metrics based on the preferences of downstream applications. It is worth noting that we do not aim to propose a new metric that supersedes all existing ones, recognizing that the best metric can depend on the specific downstream application.
@@
Given these perturbations, we can measure the sensitivity of a metric to each perturbation, and compare the sensitivity between metrics. The basic idea is that for each perturbation, we can establish an exchange rate between two metrics under this perturbation. For example, to represent the same surface orientation perturbation, we need 2 units of metric A but 1 unit of metric B. This is an exchange rate of 2:1. We can then compare exchange rates across metrics and perturbations to better understand the behavior of various metrics in terms of how they respond to different perturbations.
Note that the exchange rates must be interpreted in comparison, not in isolation. An exchange rate of 20:1 between metric A and B under a perturbation (say surface orientation) does not tell which metric is more sensitive because the unit of each metric can be chosen arbitrarily (e.g., meter instead of millimeter). But if we discover an exchange rate of 2:1 between the same two metrics using the same units under another perturbation (say relative scale), we can conclude that compared to metric B, metric A is more sensitive to surface orientation than to relative scale, because 1 unit worth of surface orientation perturbation under metric B translates to 20 units of metric A, whereas 1 unit worth of relative scale perturbation under metric B only translates to 2 units of metric A. In other words, relative to metric B, metric A amplifies the surface orientation perturbation.
Human Sensitivity: We also apply our sensitivity analysis to human judgment. Human judgment can serve as a useful reference for sensitivity comparisons because (1) many generative applications produce content for human consumption and (2) the human visual system remains the best general-purpose depth perception system and human judgment may provide clues on what is and is not important for achieving human-level visual capabilities.
@@
Using human judgment as a reference yields interesting findings about existing metrics. Humans are sensitive to affine transforms of depth or disparity, but many metrics perform affine alignment and are thus completely insensitive. Most notably, all widely used metrics have very poor sensitivity to curvature perturbation (e.g., making a flat plane wavy).
Sensitivity Aligned Composition: In addition to improving our understanding, our sensitivity analysis enables us to combine existing metrics to align with a specific sensitivity profile, such as that of humans or a downstream application. The basic idea, which we call “sensitivity aligned composition (SAC)”, is to combine existing metrics through some parametric form of composition such as a weighted average, and optimize the composition parameters such that the combined metric achieves the designed sensitivities to a given set of perturbations.
Because existing metrics are overly insensitive to curvature perturbation compared to human judgment, we propose a new metric RelNormal, which is based on relative surface normals and thus sensitive to curvature perturbation. Combining RelNormal and a subset of existing metrics, we propose SAWA-H (Sensitivity Aligned Weighted Average based on Human judgment), a new, composite metric that aligns better with human judgment than all existing metrics.
@@ 2 Related Work
Monocular Depth Evaluation Metrics. Conventional metrics directly evaluate the difference between predicted and ground truth depth. They evaluate accuracy of ordinal relationship (Zoran et al., 2015; Chen et al., 2016) or per-pixel depth difference (Eigen et al., 2014; Saxena et al., 2008). Recently, Koch et al. (2018); Chen et al. (2019); Örnek et al. (2022); Talker et al. (2024); Pham et al. (2024) noted that these metrics are not sensitive to over-smooth boundaries and error at predicting planes, and new metrics were proposed to evaluate boundary sharpness (Koch et al., 2018; Chen et al., 2019; Bochkovskii et al., 2025) and difference in 3D (Koch et al., 2018; Örnek et al., 2022; Wang et al., 2025a). Though prior work noted that existing metrics are not sensitive to some errors, there lacks a principled way to systematically study behavior of different metrics. We attempt to fill this gap.
@@
Monocular Depth Estimation Methods. There is rapid development in MDE in recent years. It is worth noting that a decent number of methods predict affine-invariant depth or disparity (Birkl et al., 2023; Yang et al., 2024a; b; Ke et al., 2024; Fu et al., 2024), and during evaluation, they perform affine alignment on depth or disparity before computing metrics. Li et al. (2025) pointed out two issues with alignment: unfair comparison between methods that use different alignment, and alignment is sensitive to outliers. In our analysis we further find that affine-aligned metrics are misaligned with human judgment, which is sensitive to geometric distortions caused by affine perturbations of depth.
@@
Human Sensitivity and Visualization Tools Though human vision is known to be accurate and fairly robust, humans still exhibit peculiarities when judging 3D structure. When constructing the OASIS dataset, Chen et al. (2020) observed that annotators judged the shape correctly (i.e. the relative normals) but often made mistakes when estimating the overall orientation. Linsley et al. (2025) find that though humans achieve similar scores as deep neural networks when estimating depth order, humans are much better at visual perspective taking (answering ”can object A see object B?”). This suggests that standard 3D geometric evaluations fail to capture the aspects of a scene that humans are sensitive to.
@@
When visualizing depth, monocular depth estimation works typically plot heatmaps of either depth or disparity (Birkl et al. (2023); Eigen et al. (2014); Zoran et al. (2015)). Other works also display novel views of the unprojected depth map with the points colored by the corresponding pixel in the original image Wang et al. (2025a); Yin et al. (2021). Notably, the shading in these point clouds does not correspond with the true geometry. Wang et al. (2025a), Cao et al. (2022), and others additionally display a gray mesh object with shadows and specular highlights rendered from the original view. This is significant, as Liu & Todd (2004) finds that shadows, specular highlights, and other aspects of normal shading greatly improve humans’ ability to discern curvature. We further expand on these visualization tools in section 4.1.
@@ 3 Sensitivity Analysis
  • Surface orientation perturbation refers to perturbing a depth map such that the surface normal of each pixel is rotated by the same amount (or close to the same amount). Fig. 1 Left shows an example.
  • Camera intrinsics perturbation refers to perturbing both focal length and depth while keeping the 3D shape similar in terms of surface normals. Fig. 1 Middle & Right shows an example: although the 3D geometry is similar, the absolute difference between predicted depth and ground truth depth is large (Fig.  1 Right).
  • Relative scale perturbation refers to perturbing relative scale between objects, e.g. scaling down the depth of a floating foreground object while keeping the background the same.
  • Curvature perturbation refers to perturbing the curvature of back-projected 3D surfaces, e.g. making a smooth plane wavy. We perform curvature perturbation under two spatial frequencies. Please refer to Appendix for details.
  • Affine transform perturbation refers to perturbing ground truth depth by an affine transform of depth or disparity.
  • Boundary perturbation refers to blurring depth values across occlusion boundaries, resulting in distortion of geometry near boundaries.
Please refer to Appendix for details of the perturbation algorithms and examples of perturbed depth.
@@
This intuitive notion of exchange rate is well defined if both metric A and B respond linearly to the intensity of the perturbation P. However, this is almost never the case. For example, some metrics are bounded and will eventually saturate for large perturbations. Therefore our formal definition of the exchange ratio needs to also handle nonlinear functions.
@@
We create a dataset of synthetic scenes using Infinigen Raistrick et al. (2023; 2024), a procedural generator of photorealistic nature and indoor scenes. We choose to use synthetic data because real-world data are limited in the diversity of scenes and availability of dense depth ground truth. For example, no pixelwise real-world ground truth is available for large natural scenes. We choose Infinigen because it covers both indoor and natural scenes and is easy to customize.
The dataset consists of 95 scenes. Among them, 56 are indoor scenes, which have short depth range and more regular shape, and 39 are nature scenes, which have long depth range and less regular shape. For each perturbation type, we perturb by at least 6 different intensities, with a total of 5320 perturbed depth in the dataset. Example scenes are included in Appendix F.
An Example of Comparison of Existing metrics Here we show an example of using our methodology to compare sensitivity between 3 metrics: AbsRel-No Align (AbsRel (Saxena et al., 2008) with no alignment), AbsRel-Disparity Af (AbsRel after affine alignment of disparity), and Boundary F1-No Align (Boundary F1 (Bochkovskii et al., 2025) with no alignment).
@@ 4 Human Sensitivity
We are particularly interested in how the sensitivity of existing metrics compares to human judgment. This is a useful question because many generative AI applications generate content for human consumption, and for such applications human judgment may be the only evaluation that matters. In addition, findings of human sensitivity can reveal what is and is not important for achieving human-level visual capabilities, which would be sufficient for robotic applications.
@@
Note that unlike computer metrics, humans are not given two depth maps to compare. Instead, humans are shown an RGB image and a (possibly) perturbed depth map. That is, humans are comparing the perturbed depth map against their own internal 3D reconstruction. This reveals human sensitivity to perturbations in the specific context of monocular depth estimation.
@@ 4.2 Sensitivity of Existing Metrics Relative to Human Judgment
Metrics: Using the dataset described in Sec. 3, we compare the sensitivity of 9 widely used metrics against human judgment: AbsRel, AbsRelp, $\delta^{0.125}$, $\delta^{1}$, RMSE, RMSE (log), RMSE (log, scale invariant), WKDR, and boundary F1. Because it is common to align prediction with ground truth before computing metrics, we compute these metrics with no alignment, or with one of the following alignments if applicable: scale alignment of depth, affine alignment of depth (L1 or least square), affine alignment of disparity, and scale or affine alignment of point map (Ke et al., 2024; Wang et al., 2025a). More details are in Appendix C.
@@
To detect low quality annotators who rate everything as “not ground truth”, we randomly select 10 scenes and set the perturbation intensity to zero. This results in annotations on a total of 315 depth maps being collected. More details are in Appendix B.
Results: Fig. 4 shows the exchange rates between every metric and human judgment under various perturbations, revealing two main findings: (1) relative to humans, existing metrics (all rows except the last four) are severely under-sensitive to the curvature perturbation; (2) humans are sensitive to affine transforms of depth or disparity, which is ignored by metrics that perform the respective affine alignment. In Figure 5 and Figure 6 we use simple shapes to illustrate the theoretical causes of these observations.

### v3

# Toward A Better Understanding of Monocular Depth Evaluation

@@ Abstract
Monocular depth estimation is an important task with rapid progress, but how to evaluate it is not fully resolved, as evidenced by a lack of standardization in existing literature and a large selection of evaluation metrics whose trade-offs and behaviors are not fully understood. This paper contributes a novel, quantitative analysis of existing metrics in terms of their sensitivity to various types of perturbations of ground truth, emphasizing comparison to human judgment. Our analysis reveals that existing metrics are severely under-sensitive to curvature perturbation such as making smooth surfaces bumpy. To remedy this, we introduce a new metric based on relative surface normals, along with new depth visualization tools and a principled method to create composite metrics with better human alignment. Code and data are available at: https://github.com/princeton-vl/evalmde.
@@ 1 Introduction
However, how to evaluate monocular depth estimation is not fully resolved. This is reflected by a lack of standardization in existing literature, and a very large menu of evaluation metrics. It is in fact possible for two papers (e.g., [bochkovskiydepthpro, wang2025moge]) to report 5 to 7 different metrics each but with zero overlap [^1] Two metrics can share the same name but are different due to differences in depth alignment..
Even when models are compared under the same set of metrics, it is often not clear how to interpret the results. If a model performs well under one metric (say AbsRel after affine alignment on disparity) but poorly under another (say $\delta^{1}$ with no alignment), what does it mean? Often there lack clear answers, because the trade-offs between different metrics are not fully understood.
The large selection of metrics arises from a combination of decisions that need to be made when designing an evaluation metric. For example, the decisions can include how to compute error of individual depth values against ground truth (by difference in depth, log depth, or inverse depth), whether and how to discount error for faraway points, whether to binarize the error and if so with what threshold, and how to account for the unknown global scale through alignment (in depth or inverse depth), and whether to also allow an unknown global shift.
The complexity of these decisions is due to several issues inherent to monocular depth estimation that preclude straightforward comparisons of individual depth values. One is scale ambiguity. The global scale of a scene is fundamentally ambiguous: a scene can be a miniature replica of a larger identical scene. This scale ambiguity can also occur locally in a scene: it can be impossible to tell from a single image whether a drone in the sky is a big one far away or small one close by. Even without any scale ambiguity, another issue is the unbounded range of possible depth values in the same scene—consider an ocean that extends to the horizon. In such cases, some form of normalization or weighting is necessary to prevent errors on faraway objects from dominating.
@@
The discussion above suggests that the evaluation of monocular depth is inherently complex. The existence of many metrics reflects best efforts to address such complexity. On the other hand, there has been limited study on the trade-offs of these metrics, and research on monocular depth could benefit from additional investigation.
In this paper we seek to improve our understanding of monocular depth evaluation. Our goal is to study various evaluation metrics and shed light on their trade-offs and behaviors, and to develop a principled method to customize or combine metrics to align with particular preferences, such as those based on human perception.
It is important to note that our goal is not a new “perfect” metric that is better in all aspects and supersedes all existing metrics. A single ideal metric probably does not exist and the best metric is likely application dependent. We only seek to offer additional insights on the behaviors of existing metrics and how they might be customized and combined toward particular goals.
@@
Given these perturbations, we can measure the sensitivity of a metric to each perturbation, and compare the sensitivity between metrics. The basic idea is to measure the numerical responses of a metric under various perturbations, and compare such responses across metrics, after proper normalization to account for the arbitrary choice of measuring units.
Human Sensitivity: We also apply our sensitivity analysis to human judgment. Human judgment can serve as a potentially useful reference for sensitivity comparisons because (1) many generative applications produce content for human consumption and (2) the human visual system remains arguably the best general-purpose depth perception system and human judgment may provide clues on what is and is not important for achieving human-level visual capabilities. Note that we do not claim that human judgment is the gold standard or an ideal proxy for all or some downstream applications; it is simply a potentially useful reference.
@@
Using human judgment as a reference yields interesting findings about existing metrics. Humans are sensitive to affine transforms of depth or disparity, but many metrics perform affine alignment and are thus completely insensitive. Most notably, all widely used metrics have very poor sensitivity to curvature perturbation (e.g., making a smooth surface bumpy).
Sensitivity Aligned Composition: In addition to improving our understanding, our sensitivity analysis enables us to combine existing metrics to align with a specific sensitivity profile, such as that of humans or a downstream application. The basic idea, which we call “sensitivity aligned composition (SAC)”, is to combine existing metrics through some parametric form of composition such as a weighted average, and optimize the composition parameters such that the combined metric achieves the desired sensitivities to a given set of perturbations.
Because existing metrics are overly insensitive to curvature perturbation compared to human judgment, we propose a new metric RelNormal, which is based on relative surface normals and thus sensitive to curvature perturbation. Combining RelNormal and a subset of existing metrics, we propose SAWA-H (Sensitivity Aligned Weighted Average based on Human judgment), a new, composite metric that aligns better with human judgment than all existing metrics. SAWA-H can supplement existing metrics in cases where alignment with human judgment is desired.
@@ 2 Related Work
Monocular Depth Evaluation Metrics. Conventional metrics directly evaluate the difference between predicted and ground truth depth. They evaluate accuracy of ordinal relationship [zoran2015ordinalrelationship, chen2016diw] or per-pixel depth difference [eigen2014depth, saxena2008make3d]. Recently, koch2018evaluation, chen2019oversmooth, ornek20222dto3d, talker2024mindedgerefiningdepth, pham2024sharpdepthsharpeningmetricdepth noted that these metrics are not sensitive to over-smooth boundaries and error at predicting planes, and new metrics were proposed to evaluate boundary sharpness [koch2018evaluation, chen2019oversmooth, bochkovskiydepthpro] and difference in 3D [koch2018evaluation, ornek20222dto3d, wang2025moge]. Though prior work noted that existing metrics are not sensitive to some errors, there lacks a principled way to systematically study behavior of different metrics. We attempt to fill this gap.
@@
Monocular Depth Estimation Methods. There is rapid development in MDE in recent years. It is worth noting that a decent number of methods predict affine-invariant depth or disparity [birkl2023midas, yang2024depthanything, yang2024depthanythingv2, ke2024marigold, fu2024geowizard], and during evaluation, they perform affine alignment on depth or disparity before computing metrics. li2025benchdepth pointed out two issues with alignment: unfair comparison between methods that use different alignment, and alignment is sensitive to outliers. In our analysis we further find that affine-aligned metrics are misaligned with human judgment, which is sensitive to geometric distortions caused by affine perturbations of depth.
@@
Human Sensitivity and Visualization Tools Though human vision is known to be accurate and fairly robust, humans still exhibit peculiarities when judging 3D structure. When constructing the OASIS dataset, chen2020oasis observed that annotators judged the shape correctly (i.e. the relative normals) but often made mistakes when estimating the overall orientation. linsley2025the find that though humans achieve similar scores as deep neural networks when estimating depth order, humans are much better at visual perspective taking (answering “can object A see object B?”). This suggests that standard 3D geometric evaluations fail to capture the aspects of a scene that humans are sensitive to.
@@
When visualizing depth, monocular depth estimation works typically plot heatmaps of either depth or disparity ([birkl2023midas, eigen2014depth, zoran2015ordinalrelationship]). Other works also display novel views of the unprojected depth map with the points colored by the corresponding pixel in the original image [wang2025moge, yin2021learning]. Notably, the shading in these point clouds does not correspond with the true geometry. wang2025moge, BiNi, and others additionally display a gray mesh object with shadows and specular highlights rendered from the original view. This is significant, as liu2004perceptual finds that shadows, specular highlights, and other aspects of normal shading greatly improve humans’ ability to discern curvature. We further expand on these visualization tools in section 4.1.
@@ 3 Sensitivity Analysis
Definition of evaluation metric Without loss of generality, we define a evaluation metric to be a function that maps a question and an answer to a non-negative score that represents the quality of the answer. For monocular depth, the question is a RGB image and the answer is a depth map. Sometimes the answer could also include predicted camera intrinsics.
This definition is general and does not assume access to a ground truth answer, because there may not be a unique correct answer. The metric can be a neural network or a human. On the other hand, an evaluation metric, like all existing monocular depth metrics, can use privileged information including a ground truth answer and return a score by comparing the answer to the ground truth.
Inherent Ambiguity in Monocular Depth: It is worth noting that monocular depth estimation has inherent ambiguity: there always exist multiple possible 3D shapes that project to the same 2D image, and some are considered more plausible only because of priors. For example, the input image can depict an actual castle or a printed photo of the same castle, and there can be no way to tell them apart.
In addition, the actual physical scene that generated the 2D image can be unreasonable as ground truth. Consider the Ames room [wikipedia:ames_room], which is trapezoidal-shaped but appears rectangular to the human eye. It would be unreasonable for the ground truth to be trapezoidal, because this would lead to a monocular depth system that reconstructs trapezoids even for actual rectangular rooms.
This inherent ambiguity suggests that (1) it is more fitting to use the more general definition of evaluation metric that does not assume a single known ground truth, and (2) human perception can be a reasonable reference “ground truth”, because the inherent ambiguity makes it impossible to always correctly recover the actual physical ground truth.
@@
  • Surface orientation perturbation refers to perturbing a depth map such that the surface normal of each pixel is rotated by the same amount (or close to the same amount). The motivation is that the absolute orientation of surfaces can be ambiguous to humans [chen2020oasis]. Fig. 1 Left shows an example.
  • Camera intrinsics perturbation refers to perturbing both focal length and depth while keeping the 3D shape similar in terms of surface normals. The motivation is that prediction of focal length can be off but the reconstructed shape can still be reasonable. Fig. 1 Middle & Right shows an example: although the 3D geometry is similar, the absolute difference between predicted depth and ground truth depth is large (Fig.  1 Right).
  • Relative scale perturbation refers to perturbing relative scale between objects, e.g. scaling down the depth of a floating foreground object while keeping the background the same. The motivation is that the scale of a (floating) foreground object can be ambiguous to determine relative to the background.
  • Curvature perturbation refers to perturbing the curvature of back-projected 3D surfaces, e.g. making a smooth surface bumpy. We perform curvature perturbation under two spatial frequencies. The motivation is the existence of smooth surfaces in nature (e.g., sand dunes, egg shells) and built environments. Please refer to the Appendix for details.
  • Affine transform perturbation refers to perturbing ground truth depth by an affine transform of depth or disparity. The motivation is that many existing methods predict affine invariant depth.
  • Boundary perturbation refers to blurring depth values across occlusion boundaries, resulting in distortion of geometry near boundaries. The motivation that some depth methods produce sharper boundaries, which are considered more desirable.
Please refer to Appendix 8 for details of the perturbation algorithms and examples of perturbed depth.
@@
This intuitive notion of exchange rate is well defined if both metric A and B respond linearly to the intensity of the perturbation P. However, this is almost never the case. For example, some metrics are bounded and will eventually saturate for large perturbations. Therefore our formal definition of the exchange rate needs to also handle nonlinear functions.
@@
We create a dataset of synthetic scenes using Infinigen [raistrick2023infinite, raistrick2024infinigenindoor], a procedural generator of photorealistic nature and indoor scenes. We choose to use synthetic data because real-world data are limited in the diversity of scenes and availability of dense depth ground truth. For example, no pixelwise real-world ground truth is available for large natural scenes. We choose Infinigen because it covers both indoor and natural scenes and is easy to customize.
The dataset consists of 95 scenes. Among them, 56 are indoor scenes, which have short depth range and more regular shape, and 39 are nature scenes, which have long depth range and less regular shape. For each perturbation type, we perturb by at least 6 different intensities, with a total of 5320 perturbed depth in the dataset. Example scenes are included in Appendix 14.
An Example of Comparison of Existing metrics Here we show an example of using our methodology to compare sensitivity between 3 metrics: AbsRel-No Align (AbsRel [saxena2008make3d] with no alignment), AbsRel-Disparity Af (AbsRel after affine alignment of disparity), and Boundary F1-No Align (Boundary F1 [bochkovskiydepthpro] with no alignment).
@@ 4 Human Sensitivity
We are particularly interested in how the sensitivity of existing metrics compares to human judgment. This is a useful question because many generative AI applications generate content for human consumption, and for such applications human judgment is highly relevant. In addition, findings of human sensitivity can reveal what is and is not important for achieving human-level visual capabilities, which might be sufficient for robotic applications.
@@
Note that unlike computer metrics, which relies on privileged access to ground truth answers, humans are not given two depth maps to compare, because two depth maps can be obviously different but indistinguishable in the terms of reconstructing the input image. Consistent with our definition of an evaluation metric, humans are shown an RGB image and a (possibly) perturbed depth map and asked to judge whether the depth map is the correct reconstruction. This reveals human sensitivity to perturbations in the specific context of monocular depth estimation.
@@ 4.2 Sensitivity of Existing Metrics Relative to Human Judgment
Metrics: Using the dataset described in Sec. 3, we compare the sensitivity of 9 widely used metrics against human judgment: AbsRel, AbsRelp, $\delta^{0.125}$, $\delta^{1}$, RMSE, RMSE (log), RMSE (log, scale invariant), WKDR, and boundary F1. Because it is common to align prediction with ground truth before computing metrics, we compute these metrics with no alignment, or with one of the following alignments if applicable: scale alignment of depth, affine alignment of depth (L1 or least square), affine alignment of disparity, and scale or affine alignment of point map [ke2024marigold, wang2025moge]. More details are in Appendix 10.
@@
To detect low quality annotators who rate everything as “not ground truth”, we randomly select 10 scenes and set the perturbation intensity to zero. This results in annotations on a total of 315 depth maps being collected. More details are in Appendix 9.
Results: Fig. 4 shows the exchange rates between every metric and human judgment under various perturbations, revealing two main findings: (1) relative to humans, existing metrics (all rows except the last four) are severely under-sensitive to the curvature perturbation; (2) humans are sensitive to affine transforms of depth or disparity, which is ignored by metrics that perform the respective affine alignment. We use simple shapes to illustrate the theoretical causes of these observations in Appendix 11.